OpenStudy (anonymous):

Which is the Standard Form of the equation y = −x − 6. x + y = −6 4x + y = −6 x + 4y = −24 −4x + y = −24

OpenStudy (anonymous):

standard form of y = mx+n is ax+by+c = 0 where m = -a/b and n = c